Determine the current through each of the specified points.

The circuit to the right consists of a battery (Vo 15.0 V) and five resistors (R1 110 R2 7820, R3 4630, R4 334 m, and Rs 3650). Find the current passing through each of the specified points Number 18.29 mA Number mA Number mA V
Answer #1

R2 , R3 , R4 , R5 are in parallel


1/R2345 = 1/R2 + 1/R3 + 1/R4 + 1/R5

1/R2345 = 1/782 + 1/463 + 1/334 + 1/365


R2345 = 109 ohm


R1 , R2345 are in series


Rtot = R1 + R2345


Rtot = 711 + 109 = 820 ohms


-------------

IE = V/Rtot = 15/820 = 18.3 mA


I2 = IE*R2345/R2 = 18.3*109/782 = 2.55 mA


I3 = V2345/R3 = IE*R2345/R3 = 18.3*109/463 = 4.32 mA

I4 = V2345/R4 = IE*R2345/R4 = 18.3*109/334 = 5.97 mA

I5 = V2345/R5 = IE*R2345/R54 = 18.3*109/365 = 5.46 mA


IB = I2 + I3 = 7.87 mA      <<<---answer

IF = I3 + I4 + I5

IF = 4.32 + 5.97 + 5.46 = 17.75 mA <<<<---answer